im trying to solve tis problem and find the least common multiple of each pair of polynomials, please help!
2x^2-18 and 5x^3+30x^2+45x

Answer by stanbon(57219)

find the least common multiple of each pair of polynomials, please help!
---
Factor each expression
2x^2-18 = 2(x^2-9) = 2(x-3)(x+3)
----------------------------
5x^3+30x^2+45x = 5x(x^2+6x+9)= 5x(x+3)^2
----
The lcm must have each of the factors in their highest power:
---
lcm = 2*5*x(x-3)(x+3)^2
============================
Cheers,
Stan H.